RETURN ON EQUITY

Commonwealth Construction (CC) needs $2 million of assets to get started, and it expects to have a basic earning power ratio of 35%. CC will own no securities, so all of its income will be operating income. If it so chooses, CC can finance up to 60% of its assets with debt, which will have an 8% interest rate. If it chooses to use debt, the firm will finance using only debt and common equity, so no preferred stock will be used. Assuming a 40% tax rate on all taxable income, what is the difference between CC's expected ROE if it finances these assets with 60% debt versus its expected ROE if it finances these assets entirely with common stock? Round your answer to two decimal places.

%
Answer #1

Basic Earning Power Ratio = EBIT/Total Assets

Thus, EBIT = Basic Earning Power Ratio * Total Assets = 35% * $2,000,000 = $700,000

If CC goes for debt funding, then the total amount of debt = 60% * $2,000,000 = $1,200,000 and the remaining 40% i.e. $800,000 will be equity.

Cost of debt = Interest rate * (1 - tax rate) = 8% * (1 - 40%) = 8% * 0.6 = 4.8%

Particulars Debt Funding Equity Funding
Debt        12,00,000 0
Equity (A)           8,00,000            20,00,000
EBIT (B)           7,00,000               7,00,000
Interest (C = 4.8% * B)              57,600 0
EBT (D = B - C)           6,42,400               7,00,000
Taxes (E = 40% *D)           2,56,960               2,80,000
PAT (F = D - E)           3,85,440               4,20,000
ROE (G = F/A) 48.18% 21.00%

The company's ROE is more when it uses debt funding compared to entire equity funding because of the tax shield that it gets for using debt as a source of finance.